A company is considering Pittsburgh, Philadelphia, and Baltimore as candidate locations for a new facility. For the weighted-factor rating technique, the company is evaluating each location on three factors (distance to customers, labor costs, and utility costs). The factors are weighted 0.5, 0.3, and 0.2, respectively. Pittsburgh received scores of 90, 75, and 65 for the three factors. Philadelphia received scores of 85, 82, and 70. Baltimore received scores of 82, 85, and 75. Which of the three cities should be selected (ranked first by the weighted-factor rating technique?
Answer #1

The total weighted factor rating score for a location is calculated as the sum of the product of the factor score and their respective weights for the particular location. So using this formula the total weighted score for each location are

  • For pittsburgh = (0.5x90)+(0.3x75)+(0.2×65) = 45+22.5+13 = 80.5
  • For Philadelphia= (0.5x85)+(0.3x82)+(0.2x70) = 42.5+24.6+14 = 81.1
  • For Baltimore = (0.5x82)+(0.3x85)+(0.2x75) = 41+25.5+15 = 81.5

So Baltimore should be selected as it has the highest weighted score of 81.5 among all the three alternatives